OpenStudy (anonymous):

okay so i cant get this equation 6(x+3) over 5=3x-3???????? im stupmed

OpenStudy (anonymous):

\[\frac{6(x+3)}{5}=3x-3\] start by multiplying both sides by 5 to get rid of the fraction

OpenStudy (anonymous):

@mamacase Can you follow the above instruction?

OpenStudy (anonymous):

im tryin to workit out

OpenStudy (anonymous):

Show your work here, pls!

OpenStudy (anonymous):

without te fractions it would b 6x+18=3x-3*5??

OpenStudy (anonymous):

close but you need to multiply 5 to the entire right side \[5*\frac{6(x+3)}{5}=(3x-3)*5\] \[6(x+3)=5(3x-3)\]

OpenStudy (anonymous):

although it looks like you understand how to distribute so im going to assume you know how to finish this problem

OpenStudy (anonymous):

sooo 6x+18=15x-15??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

im gettin 33/9 not sure how

OpenStudy (anonymous):

18+15= 33 15-6= 9 nothing is wrong with that answer just simplify the fraction

OpenStudy (anonymous):

ok so i gott anothr one 4(x-2) over 8=2(x+1) over 7? i kno 56 is te lcd!

OpenStudy (anonymous):

you dont need LCD for this you just need to cross multiply or basically just multiply the LCD to both sides

OpenStudy (anonymous):

ex. cross multiplication \[\frac{1}{3}x=\frac{1}{2}x\] \[2*x=3*x\]

OpenStudy (anonymous):

i still dnt understand